OpenStudy (anonymous):

How do you solve r+6/r-5=0 6/r is a fraction.

OpenStudy (phi):

multiply both sides of the equation by r \[ r(r+\frac{6}{r} -5)= 0 \cdot r \] you get \[ r^2 +6 -5r = 0 \] or, arranging in descending powers \[ r^2 -5r + 6 =0 \] This is a quadratic equation. In this case we can factor it: (r-2)(r-3) =0 so either r= 2 or r= 3 check: \[ 2 +\frac{6}{2} - 5 = 2+3-5= 5-5= 0 \] \[ 3 +\frac{6}{3} -5 = 3+2-5= 5-5= 0 \]